OpenStudy (anonymous):

Algebra 1 I really need help so I will give medal in the graph of the inequality x - 2y >= 4 which is a for x on the boundary line and x axis A. 4 B. -2 C. 2 D. -4

OpenStudy (anonymous):

Ok. You use the method of putting a zero on the y and x axis to see what the equal. like 0 - 2y >= 4 so -2*y =4 so y is -2 and x is 4

OpenStudy (anonymous):

To find the x value you do x-2(0)=4 so x =4

OpenStudy (anonymous):

X is 4 so its a

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

do u understand the method